OpenStudy (anonymous):

Suppose f(π/3) = 3 and f '(π/3) = −5, and let g(x) = f(x) sin x and h(x) = (cos x)/f(x). Find the following: g'(π/3)

OpenStudy (anonymous):

Use product rule: \[ g'(x) = f'(x)\sin(x)+f(x)\cos(x) \]

OpenStudy (anonymous):

Then just plug in \(x=\pi/3\).

OpenStudy (anonymous):

I get 7/2 - 5 sqrt (3/2) but this is given by the program as incorrect. Any further assistance?

OpenStudy (anonymous):

You mixed \(f'\) with \(f\)

OpenStudy (anonymous):

ah. Would the answer then be 3/2 - (5 sqrt 3)/2 ? Down to my last try so I'd rather not get it wrong. But you are right I confused the two f's before.

OpenStudy (anonymous):

\[ g'(x) = (-5)(\sqrt{3}/2)+(3)(1/2)=3/2-5\sqrt{3}/2 \]

OpenStudy (anonymous):

I'd try it.

OpenStudy (anonymous):

It was correct. Thank you very much good sir, I was stuck. I really need more sleep. You can has medal